# Theory of Microphase Separation in Elastomers

### v1

@@ I.1 Elastic deformation and material conservation
where we have introduced the displacement field $\bm{u}(\bm{x})$. Material conservation within an arbitrary subvolume $\Sigma$ before and after deformation gives [55]
$\int_{\Sigma}\mathrm{d}^{3}x\,\phi_{0}=\int_{\chi(\Sigma)}\mathrm{d}^{3}y\,\phi(\bm{y})=\int_{\Sigma}\mathrm{d}^{3}x\,\left|\det\bm{\nabla}\chi\right|\,\phi(\bm{x}).$ (S2)
In the last step in Eq. (S2), we have changed variables back to $\bm{x}$ and write $\phi(\chi(\bm{x}))$ as just $\phi(\bm{x})$ for simplicity. This introduces the Jacobian factor $\det(\bm{\nabla}\chi)$, which is always positive as physical deformations preserve orientation. Because of the arbitrariness of $\Sigma$, we equate the first and last integrands in Eq. (S2) to obtain the local material conservation relation
@@
For small deformations close to the critical point, we can expand both $\phi_{0}$ and $\phi(\bm{x})$ around $\phi=\phi_{\text{c}}$ and express $\bm{\nabla\cdot}\bm{u}$ as
$\bm{\nabla\cdot}\bm{u}=-\phi_{\text{c}}^{-1}\psi+\mathcal{O}(\psi_{0})+\mathcal{O}(\psi^{2}).$ (S4)
Here $\psi_{0}=\phi_{0}-\phi_{\text{c}}$ and $\psi(\bm{x})=\phi(\bm{x})-\phi_{\text{c}}$. This completes the derivation of Eq. (6) of the Letter.
@@ I.2 Energy of a swollen elastomer under small deformations
$\epsilon_{jk}(\bm{q})=\frac{i}{2}\left[q_{j}u_{k}(\bm{q})+q_{k}u_{j}(\bm{q})\right]\quad\text{and}\quad\bar{\epsilon}_{jk}(\bm{q})=\frac{i}{2}\left[q_{j}u_{k}(\bm{q})+q_{k}u_{j}(\bm{q})\right]K_{h}(\bm{q}),$ (S5)
@@
where $\bm{u}_{\bm{q}}^{\perp}=\bm{u}_{\bm{q}}-q^{-2}(\bm{q}\cdot\bm{u}_{\bm{q}})\bm{q}$ is the transverse component of $\bm{u}_{\bm{q}}$. After Fourier transforming the material conservation relation, Eq. (S4), which relates $\bm{u}$ and $\psi$, we can write the elastic energy (up to additive constants) in Fourier space as
@@
Here $M=\phi_{\text{c}}^{-2}(\lambda+2\mu)$ is the elastic part of the rescaled longitudinal (or the pressure-wave) modulus [33]. During phase separation, compositional changes in the polymer volume fraction arise primarily via solvent diffusion, which is not expected to affect the transverse shear modes $\bm{u}^{\perp}$. This allows us to discard them and write the total energy as
@@
$=\frac{1}{2}\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\left[a(T-T_{\text{c}})+\kappa q^{2}+MK_{h}(\bm{q})\right]\psi_{\bm{q}}\psi_{-\bm{q}}+\int\mathrm{d}^{3}x\,\left(\frac{1}{4}b\psi^{2}-\eta\psi\right).$ (S8)
In the last step in Eq. (S8), we have written $\mathscr{F}_{\text{el}}$ and the quadratic parts of both the mixing free energy $f(\psi)$ and the interfacial energy $\frac{1}{2}\kappa\left|\bm{\nabla}\psi\right|^{2}$in Fourier space. For a Gaussian kernel $K_{h}(\bm{x})=(4\pi h^{2})^{-3/2}\mathrm{e}^{-\left|\bm{x}\right|^{2}/(4h^{2})}$ we have $K_{h}(\bm{q})=\mathrm{e}^{-h^{2}q^{2}}$. From the above equation, we deduce that the Fourier transform of the effective binary interaction for the $\psi$ field is
$F_{\bm{q}}=a(T-T_{\text{c}})+\kappa q^{2}+M\text{e}^{-h^{2}q^{2}},$ (S9)
which completes the derivation of Eq. (7) of the Letter.
For computational purposes, it is also useful to consider the real-space representation of Eq. (S8), given by
$\mathscr{F}[\psi]=\int\mathrm{d}^{3}x\,\left[\frac{1}{2}a(T-T_{\text{c}})\psi^{2}+\frac{1}{4}b\psi^{4}+\frac{1}{2}\kappa\left|\bm{\nabla}\psi\right|^{2}+M\psi\bar{\psi}-\eta\psi\right],$ (S10)
@@ II Comparison with experiments
The experimental results in Ref. [24] are discussed in terms of the mass fraction $c_{\text{oil}}$ of the oil (solvent) present in the elastomer. Assuming that the mass densities of the elastomer and the oil ($\rho$ and $\rho_{\text{oil}}$, respectively) do not change considerably during the swelling process, the network volume fraction $\phi$ is estimated from $c_{\text{oil}}$ using
@@
The amount of oil absorbed depends on the dry Young’s modulus of the elastomer $Y$ and the temperature $T$. Because of this, the mass fraction $c_{\text{oil}}$ varies between $c_{\text{oil}}\approx 0.35$ ($Y=800{\kern 2.333pt}\text{kPa}$, $T=23{\kern 2.333pt}^{\circ}\text{C}$) and $c_{\text{oil}}\approx 0.80$ ($Y=10{\kern 2.333pt}\text{kPa}$, $T=80{\kern 2.333pt}^{\circ}\text{C}$). Using the density values listed in Table 1, which also lists other physical parameters, gives us a $\phi$ roughly in the range 0.2–0.7.
@@
$m_{\text{r}}$ Molecular mass of PDMS repeat unit [–$\text{Si}(\text{CH}_{3})_{2}\text{O}$–] $1.23\times 10^{-25}{\kern 2.333pt}\text{kg}$ ($74.2{\kern 2.333pt}\text{g}{\kern 2.333pt}\text{mol}^{-1}$)
@@
$\rho_{\text{oil}}$ Mass density of the solvent (heptafluorobutyl methacrylate3332,2,3,3,4,4,4-Heptafluorobutyl methacrylate (PC11102) safety data sheet (Apollo Scientific, Bredbury, UK, 2023).) $1345{\kern 2.333pt}\text{kg}{\kern 2.333pt}\text{m}^{-3}$
@@
An elastomer consists of many polymer chains cross-linked to form a polymer network, with the chains customarily assumed to follow Gaussian statistics. The elastic moduli of a swollen elastomer depends on the polymer network volume fraction $\phi$. However, the dependence is usually weak, and given that the elastomer is close to the critical point, we take $\phi\approx\phi_{\text{c}}$ for simplicity in the estimates below. Assuming that the elastomer swells isotropically from a dry state composed entirely of the polymer network, its Lamé parameters [^4] The Lamé parameter $\lambda$ being negative in Eq. (S13) is not a cause for concern because the overall stability of an elastomer is determined by both thermodynamic and elastic contributions to the free energy. Indeed, this is why it is sometimes customary to work with osmotic moduli [57, 42, 28, 58], which take into account both contributions. $\lambda$ and $\mu$ are [42, 58]
$\lambda=\nu k_{\text{B}}T\left(\frac{1}{2}\phi_{\text{c}}-\phi_{\text{c}}^{1/3}\right)\quad\text{and}\quad\mu=\nu k_{\text{B}}T\phi_{\text{c}}^{1/3}.$ (S13)
Above, $\nu$ is the strand density, defined as the average number of strands per unit volume. A strand is a segment of the polymer network between two adjacent cross-links, with no intervening cross-links in between [41]. The strand density $\nu$ is usually expressed in terms of the average molecular mass $m_{\text{s}}$ of the strands and the dry mass density of the elastomer $\rho$ as $\nu=\rho/m_{\text{s}}$ [36, 41].
Although our discussion so far has been on the elastic moduli of the swollen elastomer, the results reported in Ref. [24] are in terms of the Young’s modulus $Y$ of the dry (i.e., unswollen) elastomer. For a dry elastomer, $Y$ is given by [41]
$Y=3\nu k_{\text{B}}T=3\left(\frac{\rho}{m_{\text{s}}}\right)k_{\text{B}}T.$ (S14)
The above equation is frequently used to experimentally estimate $m_{\text{s}}$ by measuring $Y$, e.g., via dynamic mechanical analysis [59].
For swollen elastomers, it is customary to use the end-to-end distance of the strand as the characteristic size of the network [38, 39]. Assuming that the strands behave like a freely-jointed chain modified by a Flory characteristic ratio $C_{\infty}$, the root-mean-square end-to-end distance $\langle r^{2}\rangle^{1/2}$ before swelling is given by the usual expression [41]
$\langle r^{2}\rangle^{1/2}=(C_{\infty}N)^{1/2}\ell.$ (S15)
[…]
$N=\frac{m_{\text{s}}}{m_{\text{r}}}=\frac{3\rho k_{\text{B}}T}{Ym_{\text{r}}}.$ (S16)
Using the above equation, we see that for the PDMS elastomer used in the experiments, $N$ ranges from about $10^{2}$ (at $Y=800{\kern 2.333pt}\text{kPa}$) to about $10^{4}$ (at $Y=10{\kern 2.333pt}\text{kPa}$). As the elastomer is assumed to swell from a completely dry state, the distance between the cross-links increases roughly by a factor of $\phi_{\text{c}}^{-1/3}$ once it is swollen (assuming isotropy). Thus, the end-to-end distance $\xi$ between the cross-links of a swollen elastomer scales as [37, 38]
$\xi=\langle r^{2}\rangle^{1/2}\phi_{\text{c}}^{-1/3}=\phi_{\text{c}}^{-1/3}\left(\frac{3B}{Y}\right)^{1/2}\enspace\text{with}\quad B=C_{\infty}\rho\ell^{2}k_{\text{B}}T/m_{\text{r}}.$ (S17)
[…]
To estimate the microphase domain size, apart from $\xi$, we also require the rescaled longitudinal modulus $M$ of the swollen elastomer. Using Eqs. (S13) and (S14), we express $M$ in terms of the dry Young’s modulus as
$M=\phi_{\text{c}}^{-2}(\lambda+2\mu)=\phi_{\text{c}}^{-2}\left(\frac{1}{6}\phi_{\text{c}}+\frac{1}{3}\phi_{\text{c}}^{1/3}\right)Y\sim\frac{1}{3}\phi_{\text{c}}^{-5/3}Y.$ (S18)
@@
As we have remarked in the Letter, the coarse-graining length scale is $h=n\xi$, with $n$ being the average number of cross-links over which we coarse-grain in each direction. Using Eqs. (S17) and (S18) we find the dimensionless parameter $\gamma$ to be
$\gamma=Mh^{2}\kappa^{-1}=Bn^{2}\kappa^{-1}\phi_{\text{c}}^{-7/3}.$ (S19)
@@
Note that $\gamma$ is independent of the Young’s modulus $Y$, and it satisfies the condition $\gamma>1$ for the parameter values in Table 1 (even when $n=1$). We now see that the domain size $\Lambda$ scales as
@@
$\Lambda\sim 2\pi q_{\text{m}}^{-1}=2\pi h(\ln\gamma)^{-1/2}=2\pi\left[\frac{3Bn^{2}\phi_{\text{c}}^{-2/3}}{Y\ln\left(Bn^{2}\kappa^{-1}\phi_{\text{c}}^{-7/3}\right)}\right]^{1/2}.$ (S20)
This completes the derivation of Eq. (13) of the Letter.

### v2

@@ I.1 Elastic deformation and material conservation
where we have introduced the displacement field $\bm{u}(\bm{x})$ (see Fig. 1 of the Letter). Material conservation within an arbitrary subvolume $\Omega$ before and after deformation gives [43]
$\int_{\Omega}\mathrm{d}^{3}x\,\phi_{0}=\int_{\chi(\Omega)}\mathrm{d}^{3}y\,\phi(\bm{y})=\int_{\Omega}\mathrm{d}^{3}x\,\left|\det\bm{\nabla}\chi\right|\,\phi(\bm{x}).$ (S2)
In the last step of Eq. (S2), we have changed variables back to $\bm{x}$ and write $\phi(\chi(\bm{x}))$ as just $\phi(\bm{x})$ for simplicity. This introduces the Jacobian factor $\det(\bm{\nabla}\chi)$, which is always positive as physical deformations preserve orientation. Because of the arbitrariness of $\Omega$, we equate the first and last integrands in Eq. (S2) to obtain the local material conservation relation
@@
For small deformations close to the critical point, we can expand both $\phi_{0}$ and $\phi(\bm{x})$ around $\phi=\phi_{*}$ and express $\bm{\nabla\cdot}\bm{u}$ as
$\bm{\nabla\cdot}\bm{u}=-\phi_{*}^{-1}\psi+\mathcal{O}(\psi_{0})+\mathcal{O}(\psi^{2}).$ (S4)
Here $\psi_{0}=\phi_{0}-\phi_{*}$ and $\psi(\bm{x})=\phi(\bm{x})-\phi_{*}$. This completes the derivation of Eq. (6) of the Letter.
@@ I.2 Energy of a swollen elastomer under small deformations
$\varepsilon_{jk}(\bm{q})=\frac{i}{2}\left[q_{j}u_{k}(\bm{q})+q_{k}u_{j}(\bm{q})\right]\quad\text{and}\quad\bar{\varepsilon}_{jk}(\bm{q})=\frac{i}{2}\left[q_{j}u_{k}(\bm{q})+q_{k}u_{j}(\bm{q})\right]K_{h}(\bm{q}),$ (S5)
@@
where $\bm{u}_{\bm{q}}^{\perp}=\bm{u}_{\bm{q}}-(\hat{\bm{q}}\cdot\bm{u}_{\bm{q}})\hat{\bm{q}}$ is the transverse component of $\bm{u}_{\bm{q}}$ with $\hat{\bm{q}}=q^{-1}\bm{q}$. After Fourier transforming the material conservation relation, Eq. (S4), which relates $\bm{u}$ and $\psi$, we can write the elastic energy (up to additive constants) in Fourier space as
@@
Here $M=\phi_{*}^{-2}(\lambda+2\mu)$ is the elastic part of the rescaled longitudinal (or the pressure-wave) modulus [74]. During phase separation, compositional changes in the polymer volume fraction arise primarily via solvent diffusion, which is not expected to affect the transverse shear modes $\bm{u}^{\perp}$. This allows us to discard them and write the total free energy as
@@
$=\frac{1}{2}\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\left[a(T-T_{*})+\kappa q^{2}+MK_{h}(\bm{q})\right]\psi_{\bm{q}}\psi_{-\bm{q}}+\int\mathrm{d}^{3}x\,\left(\frac{1}{4}b\psi^{2}-\eta\psi\right).$ (S8)
In the last step of Eq. (S8), we have written $\mathscr{F}_{\text{el}}$ and the quadratic parts of the mixing free energy $f(\psi)$ and the interfacial energy $\frac{1}{2}\kappa\left|\bm{\nabla}\psi\right|^{2}$ in Fourier space. For a Gaussian kernel used in the Letter, $K_{h}(\bm{x})=(4\pi h^{2})^{-3/2}\mathrm{e}^{-\left|\bm{x}\right|^{2}/(4h^{2})}$, we have $K_{h}(\bm{q})=\mathrm{e}^{-h^{2}q^{2}}$. From the above equation, we deduce that the Fourier transform of the effective binary interaction for the $\psi$ field is
$F_{\bm{q}}=a(T-T_{*})+\kappa q^{2}+M\text{e}^{-h^{2}q^{2}},$ (S9)
which completes the derivation of Eq. (7) of the Letter. For computational purposes, it is also useful to consider the real-space representation of Eq. (S8), given by
$\mathscr{F}[\psi]=\int\mathrm{d}^{3}x\,\left[\frac{1}{2}a(T-T_{*})\psi^{2}+\frac{1}{4}b\psi^{4}+\frac{1}{2}\kappa\left|\bm{\nabla}\psi\right|^{2}+\frac{1}{2}M\psi\bar{\psi}-\eta\psi\right],$ (S10)
@@ II Comparison with experiments
II.1 Rubber elasticity
Elastomers are broadly defined as crosslinked polymers that display rubber-like elasticity—a property that primarily arises from changes in the configurational entropy of the chains between the crosslinks in the polymer network [37, 38, 39]. Classical theories of entropic elasticity, however, are often insufficient to explain experimental data, and several theories, with varying levels of sophistication, have been devised to describe rubbery materials [44]. For instance, the behavior of polymer networks can significantly change if the polymer is crosslinked in the melt state, which can introduce entanglement effects that cannot be described using classical theories. In comparison, classical theories are often sufficient to describe polymer networks crosslinked in semi-dilute solutions [69]. PDMS elastomers, such as the ones we consider in this work, can be particularly susceptible to chain entanglement as they are typically crosslinked in the melt state.
To understand the effect of entanglement interactions, we consider the localization model introduced and refined by Douglas, Gaylord, Horkay, and coworkers [79, 80, 81, 82, 44, 45]. In this model, the nonlinear hyperelastic energy density $\mathscr{W}$, which is a function of the principal stretches $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$, is taken to be of the form
$\mathscr{W}(\alpha_{1},\alpha_{2},\alpha_{3})=\underbrace{\frac{1}{2}\left(1-\frac{2}{\vartheta}\right)\nu k_{\text{B}}T\left(\alpha_{1}^{2}+\alpha_{2}^{2}+\alpha_{3}^{3}-3\right)}_{\text{network elasticity}}+\underbrace{\left[G_{\text{N}}+\left(1-\frac{2}{\vartheta}\right)\omega\nu k_{\text{B}}T\right]\left(\alpha_{1}+\alpha_{2}+\alpha_{3}-3\right)}_{\text{entanglement effects}}.$ (S12)
Here $\nu$ is the strand density, defined as the average number of strands per unit volume. A strand is a segment of the polymer network between two adjacent crosslinks, with no intervening crosslinks in between. Also, $\vartheta$ is the functionality (branch number) of the polymer network. The first term in Eq. (S12) is the usual free energy density for network elasticity that is proportional to the strand density $\nu$. The second term captures the free energy changes due to chain entanglement. In this term, $G_{\text{N}}$ is the plateau modulus of the polymer melt, which can be interpreted as the shear modulus of the polymer in the near absence of crosslinks (i.e., when $\nu\to 0$). Additionally, there is also an elastic energy contribution due to entanglements that scales in proportion to the strand density $\nu$ with a material-dependent prefactor $\omega$.
Assuming an incompressible dry (unswollen) polymer network under a uniaxial stretch with $\alpha_{1}=\alpha$ and $\alpha_{2}=\alpha_{3}=\alpha^{-1/2}$, the tensile stress $\sigma_{\text{t}}$, which is what one measures in indentation tests, can be obtained from Eq. (S12) as [39]
$\sigma_{\text{t}}=\alpha\frac{\partial\mathscr{W}}{\partial\alpha}=\left[3\left(1-\frac{2}{\vartheta}\right)\nu k_{\text{B}}T\left(1+\frac{1}{2}\omega\right)+\frac{3}{2}G_{\text{N}}\right](\alpha-1)+\mathcal{O}[(\alpha-1)^{2}].$ (S13)
Noting that the linear strain $\approx\alpha-1$, we can read off the (dry) Young’s modulus of the elastomer to be
$Y=3\left(1-\frac{2}{\vartheta}\right)\nu k_{\text{B}}T\left(1+\frac{1}{2}\omega\right)+\frac{3}{2}G_{\text{N}}.$ (S14)
Figure S1: Young’s modulus $Y$ of the elastomers considered in Ref. [24] measured in indentation tests as a function of the crosslinker concentration. The strand density $\nu$ and the crosslinker concentration have a monotonic relationship, and we see no visible plateau modulus $G_{\text{N}}$ as $\nu\to 0$, showing that entanglement contribution to network elasticity is minimal.
We expect the number of crosslinks to monotonically increase with the crosslinker concentration, which we take to be a proxy for the strand density $\nu$. This assumption, for instance, was used to estimate $\nu$ for natural rubber based on the premise that each crosslinker molecule decomposes to form one tetrafunctional crosslink [46, 45]. The value of the plateau modulus $G_{\text{N}}$ can be estimated from Eq. (S14) via indentation tests in the limit $\nu\to 0$. From previous studies on polymer networks crosslinked in the melt, $G_{\text{N}}$ values are expected to be a few hundred kPa, e.g., for natural rubber $G_{\text{N}}\sim 560{\kern 2.333pt}\text{kPa}$ [45] and for PDMS, $G_{\text{N}}\sim 240\text{--}250{\kern 2.333pt}\text{kPa}$ [82].
Figure S1 shows the measured Young’s moduli for the elastomers in Ref. [24] as a function of the crosslinker concentration. Remarkably, the data indicates no significant nonzero intercept for small $\nu$. Additionally, the smallest measured value of $Y=10{\kern 2.333pt}\text{kPa}$, is substantially lower than the expected range for $G_{\text{N}}$. This strongly suggests that entanglement effects are negligible in these elastomers. Given the absence of more detailed and systematic indentation tests, at the moment, we can only attribute the lack of a measurable $G_{\text{N}}$ to the specific crosslinking technique used in Ref. [24]. In the absence of experimental estimates for the parameter $\omega$ in Eq. (S12), the choice $\omega=1/3$ has been suggested [81]. However, such a choice will not alter our numerical results significantly and because of the minimal evidence for entanglement effects, from here on, we take both $G_{\text{N}}$ and $\omega$ to be zero. This leads to the classical James–Guth “phantom” model of network elasticity [47, 48, 38], with the Young’s modulus of the elastomer in the dry state being
$Y=3\left(1-\frac{2}{\vartheta}\right)\nu k_{\text{B}}T=3\left(1-\frac{2}{\vartheta}\right)\frac{\rho k_{\text{B}}T}{m_{\text{s}}}.$ (S15)
Here we have expressed the strand density $\nu$ in terms of the average molecular mass $m_{\text{s}}$ of the strands and the dry mass density of the elastomer $\rho$ as $\nu=\rho/m_{\text{s}}$ [39]. The above equation is frequently used to experimentally estimate $m_{\text{s}}$ by measuring $Y$ [49].
The experimental results in Ref. [24] are discussed in terms of the Young’s modulus of the dry elastomer, given by Eq. (S15). However, the rescaled longitudinal modulus $M$ that enters the free energy in Eq. (S10) is that of the swollen elastomer. Swelling can change the elastic moduli of elastomers significantly, with the moduli acquiring a dependence on the polymer network volume fraction $\phi$ [82]. We assume that the elastomer swells isotropically from a dry state composed entirely of the polymer network. The elastic response of a swollen polymer network is also crucially affected by chain entanglement. As we judge chain entanglement to be minimal for the elastomers we consider, we shall continue to use results from classical rubber elasticity. By linearizing the Cauchy stress derived from Eq. (S12) for infinitesimal displacements superimposed on top of the initial swelling, we find the associated Lamé parameters [^2] The Lamé parameter $\lambda$ being negative in Eq. (S16) is not a cause for concern because the overall stability of a swollen elastomer is determined by both thermodynamic and elastic contributions to the free energy. Indeed, this is why it is sometimes customary to work with osmotic moduli [50, 28, 41, 51], which take into account both contributions. See Sec. IV for more details. $\lambda$ and $\mu$ to be [28, 51]
$\lambda=-\left(1-\frac{2}{\vartheta}\right)\nu k_{\text{B}}T\phi^{1/3}\quad\text{and}\quad\mu=\left(1-\frac{2}{\vartheta}\right)\nu k_{\text{B}}T\phi^{1/3}.$ (S16)
@@ II.1 Rubber elasticity
We can estimate the polymer volume fraction $\phi$ from the reported mass fraction $c_{\text{oil}}$ of the oil (solvent) present in the swollen elastomer. Assuming that the mass densities of the elastomer and the oil ($\rho$ and $\rho_{\text{oil}}$, respectively) do not change considerably during the swelling process, the network volume fraction $\phi$ is estimated from $c_{\text{oil}}$ using
@@
The amount of oil absorbed depends on the dry Young’s modulus of the elastomer $Y$ and the temperature $T$. Because of this, the mass fraction $c_{\text{oil}}$ reported in Ref. [24] varies between $c_{\text{oil}}\approx 0.35$ ($Y=800{\kern 2.333pt}\text{kPa}$, $T=23{\kern 2.333pt}^{\circ}\text{C}$) and $c_{\text{oil}}\approx 0.80$ ($Y=10{\kern 2.333pt}\text{kPa}$, $T=80{\kern 2.333pt}^{\circ}\text{C}$). The density values listed in Table 1, which also lists other physical parameters, gives us a $\phi$ roughly in the range 0.2–0.7. In this range, the moduli $\lambda$ and $\mu$ only have a weak dependence on $\phi$, and given that the system is close to the critical point, we take $\phi\approx\phi_{*}$ for simplicity in all our estimates. Using Eqs. (S15) and (S16), we can finally express the rescaled longitudinal modulus $M$ in terms of the dry Young’s modulus as
$M=\phi_{*}^{-2}(\lambda+2\mu)=\frac{1}{3}\phi_{*}^{-5/3}Y.$ (S18)
II.2 Choice of the length scale h and predicted domain size
In elastomers and gels, the end-to-end distance of the strands within the polymer network is typically used as the characteristic size of the network [85, 86]. Assuming that the strands behave like a freely-jointed chain modified by a Flory characteristic ratio $C_{\infty}$, the root-mean-square end-to-end distance $\xi$ before swelling is given by the usual expression [37, 38, 39]
$\xi^{2}\sim\left(1-\frac{2}{\vartheta}\right)C_{\infty}N\ell^{2}.$ (S19)
[…]
$N=\frac{m_{\text{s}}}{m_{\text{r}}}=\frac{3\rho\,k_{\text{B}}T}{2Ym_{\text{r}}}.$ (S20)
Using the above equation, we see that for the PDMS elastomer used in the experiments, $N$ ranges from about $60$ (at $Y=800{\kern 2.333pt}\text{kPa}$) to about $5000$ (at $Y=10{\kern 2.333pt}\text{kPa}$). Thus, the end-to-end distance $\xi$ between the crosslinks of an elastomer scales as [84, 85]
$\xi\sim\left(\frac{3B}{Y}\right)^{1/2}\enspace\text{with}\quad B=\frac{C_{\infty}\rho\ell^{2}k_{\text{B}}T}{4m_{\text{r}}}.$ (S21)
[…]
@@ II.2 Choice of the length scale h and predicted domain size
$m_{\text{r}}$ Molecular mass of PDMS repeat unit $\text{Si}(\text{CH}_{3})_{2}\text{O}$–] $1.23\times 10^{-25}{\kern 2.333pt}\text{kg}$ ($74.2{\kern 2.333pt}\text{g}{\kern 2.333pt}\text{mol}^{-1}$)
@@
$\rho_{\text{oil}}$ Mass density of the solvent (heptafluorobutyl methacrylate4442,2,3,3,4,4,4-Heptafluorobutyl methacrylate (PC11102) safety data sheet (Apollo Scientific, Bredbury, UK, 2023).) $1345{\kern 2.333pt}\text{kg}{\kern 2.333pt}\text{m}^{-3}$
@@
$T_{*}$ Critical temperature in the Landau free energy $f(\psi)$ [Eq. (5)] $70{\kern 2.333pt}{{}^{\circ}}\text{C}$
$\phi_{*}$ Critical volume fraction in $f(\psi)$ 0.2
$a$ Quadratic coefficient in $f(\psi)$ $0.025{\kern 2.333pt}\text{kPa}{\kern 2.333pt}\text{K}^{-1}$
$b$ Quartic coefficient in $f(\psi)$ $2{\kern 2.333pt}\text{kPa}{\kern 2.333pt}\text{K}^{-1}$
$n$ Average number of crosslinks coarse-grained over 110
@@
As we have remarked in the Letter, the coarse-graining length scale is $h=n\xi$, with $n$ being the average number of crosslinks over which we coarse-grain in each direction. Using Eqs. (S21) and (S18) we find the dimensionless parameter $\gamma$ to be
$\gamma=Mh^{2}\kappa^{-1}=Bn^{2}\kappa^{-1}\phi_{*}^{-5/3}.$ (S22)
@@
Note that $\gamma$ is independent of the Young’s modulus $Y$, and it satisfies the condition $\gamma>1$ for the parameter values in Table 1. We now see that the domain size $\Lambda$ scales as
@@
$\Lambda\sim 2\pi q_{\text{m}}^{-1}=2\pi h(\ln\gamma)^{-1/2}=2\pi\left[\frac{3Bn^{2}}{Y\ln\left(Bn^{2}\kappa^{-1}\phi_{*}^{-5/3}\right)}\right]^{1/2}=2\pi\left[N\frac{C_{\infty}n^{2}\ell^{2}}{\ln\left(Bn^{2}\kappa^{-1}\phi_{*}^{-5/3}\right)}\right]^{1/2}\hskip-10.00002pt.$ (S23)
This completes the derivation of Eq. (12) of the Letter. In the last step above, using Eq. (S20), we have expressed $\Lambda$ in terms of the number $N$ of repeat units between the crosslinks of the polymer network. A similar result was predicted by de Gennes [100] for the microphase domain size in crosslinked polymer blends using a phenomenological model different from our nonlocal model.
Figure S2: (a) General phase diagram in the $(\phi_{0},T)$ plane obtained by choosing general, unitless parameters $T_{*}=\phi_{*}=0$, $a=b=\kappa=h=1$, and $M=\mathrm{e}$ (Euler’s number). With these choices, the order parameter $\psi=\phi$, $\gamma=Mh^{2}/\kappa=\mathrm{e}$, and $q_{\text{m}}^{2}=h^{-2}\ln\gamma=1$, giving rise to domains of size $\Lambda\sim 2\pi$. Above the temperature $T_{*}^{\prime}=T_{\text{micro}}(0)=-2$ only the uniform (U) phase is stable. Microphase separation occurs below $T_{*}^{\prime}$ and leads to the formation of phases composed of stripes (S), droplets (D), and inverted droplets (ID). Close to $T_{*}^{\prime}$, there are also four regions of phase coexistence (shaded in blue) between the uniform phase and inverted droplets (U + ID), between inverted droplets and stripes (ID + S), between stripes and droplets (S + D), and between droplets and the uniform phase (D + U). (b)–(e) Equilibrium profiles at T=2.5T=-2.5 for various values of the mean order parameter $\phi_{0}$ (marked in the phase diagram). Solvent-rich ($\phi<0$) and solvent-deficient ($\phi>0$) regions are highlighted in red and blue, respectively. The equilibrium profiles were obtained by numerically minimizing the free energy (Sec. VI) in a square domain of size $\times36$.